Pr(A) = 0.36, Pr(B) = 0.3. Pr(A|B) = 0.46. What is Pr(~A and B) (probability of not A, and B)?
Answer #1

Pr(A|B) = 0.46
Pr(B) = 0.30

Pr(A|B) = Pr(A and B) / Pr(B)
0.46 = Pr(A and B) / 0.30
Pr(A and B) = 0.138

Pr(not A and B) = Pr(B) - Pr(A and B)
Pr(not A and B) = 0.30 - 0.138
Pr(not A and B) = 0.162